HUBBLE SPACE TELESCOPE DAILY REPORT #1503 PERIOD COVERED: 0000Z (UTC) 10/24/95 - 0000Z (UTC) 10/25/95 Daily Status Report as of 298/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed WF/PC-2 6119 (Snapshot Survey for Companions of Planetary = Nebula Nuclei) The WF/PC-2 was used to make snapshot survey observations of = planetary nebula NGC-2371 to study its nuclei and to look for possible = companions. The observations were executed as scheduled, and no anomalies = were noted. 1.2 Completed WF/PC-2 6187 (WF/PC-2 Cycle 5 Calibration: Earth Flats) The WF/PC-2 was used to make Earth streak flat calibration images. = These streak flats will be used to produce flat fields, to map the OTA = illumination pattern and, in combination with other data, to produce new = superflats. The observations were completed as planned, and no problems = were reported. 1.3 Completed Three Sets of WF/PC-2 6188 (WF/PC-2 Cycle 5 Calibration: = Dark) The WF/PC-2 was used to make internal dark calibration images. The = observations were completed as scheduled, and no problems were reported. 1.4 Completed FOS/RD 6019 (Supersonic Velocity Dispersion in Dwarf Nova = Accretion Disks) The Faint Object Spectrograph (Red Side) was used to make = spectrographic observations of the dwarf nova X-Cha (aka, OY Carat in = order to study its supersonic velocity dispersion. The intent is to nail = down the physical support mechanism for its accretion disk of this short = period eclipsing dwarf nova. The observations were executed as scheduled, = and no anomalies were noted. 1.5 Completed Three Sets of WF/PC-2 6253 (Pure Parallel Archive = Program) The WF/PC-2 was used to execute a pure parallel program taking = ultraviolet observations running concurrently with prime science = observations. The intent is to search for extra-galactic sources with = flux excesses at 3000 A0 and galactic sources with excesses at 1600 A0. = The observations were completed as planned, and no problems were reported. 1.6 Completed FOS/BL 5346 (B-Supergiants in Post MS Gap) The FOS (Blue Side) was utilized to look at the predicted gap in = the H-R diagram of SK-68-11 to see if it is filled with early B = supergiants. The instrument will obtain high S/N spectral data of four = early B supergiants in the LMC to discriminate between various = evolutionary scenarios and to constrain new stellar evolution scenarios. = There were no reported operational problems. 1.7 Completed WF/PC-2 6189 (WF/PC-2 Cycle 5 Calibration: Internal = Monitor - Vis Flats) The WF/PC-2 was used to make internal visual flat calibration = images. Both photometric and narrow-band filters were used. The = observations were executed as planned, and no problems were reported. 1.8 Completed Three Sets of FOS/BL 6166 (FOS Cycle 5: Location of = Spectra and Wavelength Calibration) The Faint Object Spectrograph (Blue Side) was used to make = wavelength calibration images to measure the stability of the FOS Blue = Side wavelength scale and spectral locations. The observations were = completed as planned, and no anomalies were reported. 1.9 Completed GHRS/2 5961 (The D/H Ratio in Double QSO Absorption = System Towards Q0454+220) The High Resolution Spectrograph (Side Two) was used to make = spectrographic observations of the Lyman Alpha line toward QSO Q0454-220 = in order to study the Deuterium to Hydrogen ratio. A wavelength = calibration image was also taken. The observations were completed as = planned, and no problems were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 4 Successful : 4 Scheduled Reacquisitions : 10 Successful : 10 2.2 FHST Updates: Scheduled : 10 Successful : 10 2.3 Operations Notes: There were data inversions during STGT operations due to a command = hit on TDW. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH -END-